OpenStudy (anonymous):

When an airplane leaves the runway, its angle of climb is 19° and its speed is 205 feet per second. How long will it take the plane to climb to an altitude of 10,000 feet? 16,000 feet?

OpenStudy (jdoe0001):

well, the plane is climbing at 19 degrees angle so say we'd like to know how long it'd take when it reaches 10,000 feet so firstly... we may want to find how long horizontally has it travelled by the time it reaches 10,000 up so  now recall your SOH CAH TOA so we're after the "adjacent side", but we only know the angle and opposite thus we'd use the tangent identity thus \(\bf tan(19^o)=\cfrac{opposite}{adjacent}\to \cfrac{y}{x}\implies tan(19^o)=\cfrac{10,000}{x}\implies x=\cfrac{10,000}{tan(19^o)}\)

OpenStudy (jdoe0001):

hmm anyhow \(\bf tan(19^o)=\cfrac{opposite}{adjacent}\to \cfrac{y}{x}\implies tan(19^o)=\cfrac{10,000}{x} \\ \quad \\ \implies x=\cfrac{10,000}{tan(19^o)}\)

OpenStudy (anonymous):

So know solve for which is 29042?

OpenStudy (anonymous):

solve for "x"*

OpenStudy (jdoe0001):

we know is going 205 feet per second so if we divide the horizontal by 205, should give us how many seconds it took to cover that horizontal distance yeap

OpenStudy (jdoe0001):

29042 yeap so 20942/205 thus 141.668 or 142 seconds

OpenStudy (jdoe0001):

so to climb 10,000 feet UP it also travelled 29042 feet horizontally and it covered that in 142 seconds or 2mins and 22 secs

OpenStudy (anonymous):

WOW!!

OpenStudy (anonymous):

THANK-YOU

OpenStudy (jdoe0001):

and to find how long for 16,000 just use that as the opposite side 

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

I didnt know what in the heck to do with the 205!